Skatercat:

A train leaves the station at 9:00 pm traveling east at 36 miles per hour. A second train leaves the station at 10:00 pm traveling west at 42 miles per hour. At what time will both trains have traveled the same distance? A. 4:00 am B. 5:00 am C. 6:00 am D. 7:00 am

mhchen:

So first train leaves 1 hour before the second train. Let x be hours after 10pm First train distance = 36x + 36 Second train distance = 42x Right? Now set them equal to each other 36x+ 36 = 42x Get it so far?

Skatercat:

yeah but what time will both trains get to the station if they traveled the same distance

mhchen:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Skatercat but what time will both trains get to the station if they traveled the same distance \(\color{#0cbb34}{\text{End of Quote}}\) What time will the trains get to the station? I thought they're trying to leave it.

Skatercat:

thats WHAT IM GETTING AT

Skatercat:

At what time will both trains have traveled the same distance? A. 4:00 am B. 5:00 am C. 6:00 am D. 7:00 am

Skatercat:

ummmm never mind oops -.-;

mhchen:

So you can first solve for x 36x+ 36 = 42x And that'll be the number of hours after 10pm which they'll be at the same distance. Then you can add it to 10pm

Skatercat:

ooooooooohhhhhhh